# A Constant Factor Approximation for Directed Feedback Vertex Set in Graphs of Bounded Genus

### v3

@@ 9                  $S\leftarrow S\backslash\{s_{t}\}$
16Use 1 to find FVS $\tilde{S}$ of $G^{\bar{S}}$ of cost at most $O(g)OPT_{LP}$
return $S\cup\tilde{S}$
@@
Clearly, $\bar{S}\cup\tilde{S}$ is a feasible FVS, we claim it has cost $O(g)OPT_{LP}$. It suffices to prove that $\bar{S}=O(g)OPT_{LP}$. To do so we apply that standard analysis of primal-dual methods in [11, 12].
@@
[…]
@@
Denote $\mathcal{C}^{\prime}$ the set of facial cycles of $\mathcal{C}$. For a node $v\in\bar{S}$, call a cycle $C\in\mathcal{C}^{\prime}$ with $C\cap\bar{S}$ a witness for $v$. Since we applied reverse deletion to $\bar{S}$ at the end of the algorithm, each node of $\bar{S}$ has a witness in $\mathcal{C}^{\prime}$ which is a cycle of $G^{S}$.
@@
[12] There exists a laminar family $\mathcal{A}\subset\mathcal{C}^{\prime}$ of witness cycles in $G^{S_{\bar{t}}}$ for $\bar{S}$.
@@
For the purpose of bounding $\sum_{C\in\mathcal{C}_{t}}|\bar{S}\cap C|$, we define the following bipartite graph.
@@
Since each $C\in\mathcal{C}_{t}$ is incident to the vertices of $\bar{S}$ on $C$, $|\bar{S}\cap C|$ is the degree of $C$ in $\mathcal{D}_{G}$. Summing this equality over each $C\in\mathcal{C}_{t}$ yields $\sum_{C\in\mathcal{C}_{t}}|\bar{S}\cap C|=E(\mathcal{D}_{G})$. By placing the node of the debit graph corresponding to $C$ inside the inside region of $C$ we can see that the debit graph is also embedded on $Q$.
@@
Euler’s formula (for instance see [16]) for graphs embedded on a surface of genus $g$ yields $2-2g=|V(\bar{G})|-|E(\bar{G})|+|F(\bar{G})|$. Following the same method as the proof of Euler’s formula for bipartite planar graphs with at least 3 vertices, (for instance see Corollary 4.2.10 of [6]) we observe that each face of $\bar{G}$ having at least 4 edges means $|F(\bar{G})|\leq\frac{1}{2}||$. Thus for $|V(\bar{G})|\geq 3$, $|E(\bar{G})|\leq 2|V(\bar{G})|-4+4g\leq(2+g)|V(\bar{G})|-4$. If $|V(\bar{G})|\leq 2$ then $|E(\bar{G})|\leq 1\leq(2+g)|V(\bar{G})|-3$.
For a node $w$ of $T$ that is not a leaf or the root, the subgraph of $\mathcal{D}_{G}$ induced by $\mathcal{C}_{w}\cup\bar{S}_{w}$ is embedded on $Q$ and further $|\mathcal{C}_{w}\cup\bar{S}_{w}|\geq 3$, thus by 10,
@@
This shows that $\bar{S}$ has cost $O(g)OPT_{LP}$ and hence our algorithm returns a solution of cost $O(g)OPT_{LP}$. ∎
Appendix A Statement and proofs of topological results we use
First let us prove 7.
@@ Appendix A Statement and proofs of topological results we use
We use the following theorem.

### v4

@@ 9                  $S\leftarrow S\backslash\{s_{t}\}$
return $S$
@@
Clearly, $\bar{S}$ is a feasible hitting set for the set of facial dicycles of $G$, we claim it has cost $O(g)OPT_{LP}$. To do so we apply that standard analysis of primal-dual methods in [11, 12].
@@
([11]): Suppose $S\subset V(G)$ and $y$ is a solution to ($\mbox{D}_{DFVS}$) output by our primal-dual algorithm such that the following holds.
@@
Again we remove nodes of $G$ that do not lie on any dicycle. Denote $\bar{S}_{t}$ to be the nodes of $\bar{S}$ that intersect a cycle of $\mathcal{C}_{t}$. So it suffices to show $\sum_{C\in\mathcal{C}_{t}}|\bar{S}_{t}\cap C|\leq O(g)|\mathcal{C}_{t}|.$
@@
Denote $\mathcal{C}^{\prime}$ the set of facial cycles of $\mathcal{C}$. For a node $v\in\bar{S}$, call a cycle $C\in\mathcal{C}^{\prime}$ with $C\cap\bar{S}=\{v\}$ a witness for $v$. Since we applied reverse deletion to $\bar{S}$ at the end of the algorithm, each node of $\bar{S}$ has a witness in $\mathcal{C}^{\prime}$ which is a cycle of $G^{S}$.
@@
[12] There exists a laminar family $\mathcal{A}\subset\mathcal{C}^{\prime}$ of witness cycles in $G^{S_{\bar{t}}}$ for $\bar{S}_{t}$.
@@
To bound $\sum_{C\in\mathcal{C}_{t}}|\bar{S}_{t}\cap C|$, we define the following bipartite graph.
@@
Since each $C\in\mathcal{C}_{t}$ is incident to the vertices of $\bar{S}_{t}$ on $C$, $|\bar{S}_{t}\cap C|$ is the degree of $C$ in $\mathcal{D}_{G}$. Summing this equality over each $C\in\mathcal{C}_{t}$ yields $\sum_{C\in\mathcal{C}_{t}}|\bar{S}\cap C|=E(\mathcal{D}_{G})$. By placing the node of the debit graph corresponding to $C$ inside the inside region of $C$ we can see that the debit graph is also embedded on $Q$.
@@
Euler’s formula (for instance see [17]) for graphs embedded on a surface of genus $g$ yields $2-2g=|V(\bar{G})|-|E(\bar{G})|+|F(\bar{G})|$. Following the same method as the proof of Euler’s formula for bipartite planar graphs with at least 3 vertices, (for instance see Corollary 4.2.10 of [6]) we observe that each face of $\bar{G}$ having at least 4 edges means $|F(\bar{G})|\leq\frac{1}{2}||$. Thus, for $|V(\bar{G})|\geq 3$, $|E(\bar{G})|\leq 2|V(\bar{G})|-4+4g\leq(2+g)|V(\bar{G})|-4$. If $|V(\bar{G})|\leq 2$ then $|E(\bar{G})|\leq 1\leq(2+g)|V(\bar{G})|-3$.
For a node $w$ of $T$ that is not a leaf or the root, the subgraph of $\mathcal{D}_{G}$ induced by $\mathcal{C}_{w}\cup\bar{S}_{w}$ is embedded on $Q$ and further $|\mathcal{C}_{w}\cup\bar{S}_{w}|\geq 3$, thus by 4,
@@
This shows that $\bar{S}$ has cost $O(g)OPT_{LP}$ and hence our algorithm returns a solution of cost $O(g)OPT_{LP}$.
Now let us show that in the case $G$ is embedded in a way such that there exist regions $R_{1},R_{2}$ of $Q$ homeomorphic to the open disk, such that the inside region of any facial dicycle contains at least one of $R_{1},R_{2}$, then algorithm 1 is an $8$- approximation.
The proof works exactly the same as the general case. The key here is to note that the inside regions of face minimal dicycles do not intersect. Thus, $R_{1}$ lies in the inside region of at most one cycle in $\mathcal{C}_{t}$. Likewise, $R_{2}$ lies in the inside region of at most one cycle in $\mathcal{C}_{t}$. Since inside region of any facial dicycle contains at least one of $R_{1},R_{2}$, $|\mathcal{C}_{t}|\leq 2$. Again 1 holds. For a facial dicycle $A$, denote $\operatorname{inside}_{\mathcal{C}_{t}}(A)$ the set of cycles of $\mathcal{C}_{t}$ that lie in the closure of the inside region of $A$.
Lemma 2.
There do not exist distinct $A_{1},A_{2},A_{3}\in\mathcal{A}$ such that $\operatorname{inside}_{\mathcal{C}_{t}}(A_{1})=\operatorname{inside}_{\mathcal{C}_{t}}(A_{2})=\operatorname{inside}_{\mathcal{C}_{t}}(A_{3})$.
@@
Suppose such $A_{1},A_{2},A_{3}$ existed. Since they are laminar we may assume w.l.o.g that $A_{1}$ is contained in the closure of the inside region of $A_{2}$ and $A_{2}$ is contained in the closure of the inside region of $A_{3}$. Let $v_{i}$ be the hit node that $A_{i}$ is the witness of. Note that $v_{2}$ does not lie on $A_{1}$. Thus, as $v_{2}$ lies outside $\operatorname{inside}(A_{1})$, it lies outside the closure $\operatorname{cl}(\operatorname{inside}(A_{1}))$ of $\operatorname{inside}(A_{1})$. So $v_{2}$ lies in $Q\backslash\operatorname{inside}(A_{3})$. Thus, $v_{2}$ does not lie on any cycle of $\operatorname{inside}_{\mathcal{C}_{t}}(A_{3})$. Also $v_{2}$ does not lie on $A_{3}$. Thus, as $v_{2}$ lies inside $\operatorname{cl}(\operatorname{inside}(A_{3}))$, it lies inside $\operatorname{inside}(A_{3})$. Thus, $v_{2}$ does not lie on any cycle of $\mathcal{C}_{t}\backslash\operatorname{inside}_{\mathcal{C}_{t}}(A_{3})$.
This implies that $v_{2}$ does not lie on any cycle of $\mathcal{C}_{t}$, which is a contradiction.
This implies that $|\bar{S}_{t}|=|\mathcal{A}|\leq 2(2^{|\mathcal{C}_{t}|})\leq 8$. Thus, $\sum_{C\in\mathcal{C}_{t}}|\bar{S}_{t}\cap C|\leq|\bar{S}_{t}||\mathcal{C}_{t}|\leq 8|\mathcal{C}_{t}|$. This shows algorithm 1 is an 8-approximation. ∎
4 Solving the case of no facial cycles
We now show the LP gap of the natural LP ($\mbox{P}_{\text{DFVS}}$) for $G$ has integrality gap $O(g)$ in the case $G$ contains no facial cycles. This will allow us to derive an $O(g)$-approximation for the general case by first using Theorem 2 to obtain a hitting set $S$ for the set of facial cycles of cost at most $O(g)OPT$ and then obtaining a hitting set $\bar{S}$ for the remaining dicycles.
Lemma 3.
Suppose $G$ is a digraph embedded on a surface $Q$ of some fixed genus $g$ and there is no facial dicycle of $G$. Then the LP gap of the natural LP ($\mbox{P}_{\text{DFVS}}$) for $G$ has integrality gap $O(g)$.
Proof.
We prove the statement by induction on the genus $g$. The case $g=0$ is trivial because all cycles in planar graphs are facial. Suppose the statement is true for $g=g^{\prime}$. Let $Q$ be a surface of genus $g$, Let $G$ be a digraph embedded on $Q$.
First, while the optimal solution $\bar{x}$ to ($\mbox{P}_{\text{DFVS}}$) has a vertex $v$ with value $\bar{x}_{v}\geq\frac{1}{24}$ add v to our temporary hitting set $F$. Formally initialize $F=\emptyset$. While the optimal solution $\bar{x}$ to ($\mbox{P}_{\text{DFVS}}$) for $G^{F}$ contains a value $\bar{x}_{v}$ which is $\frac{1}{24}$ or more add $v$ to $F$.
We now seek to define (integral) distances on $G^{F}$. By standard LP theory, $\hat{x}$ has rational coordinates. Let $N\in\mathbb{Z}_{>0}$ be such that $N\hat{x}$ and $\frac{1}{12}N$ are integral, call $N\hat{x}_{v}$, the weight of $v$. Define the weighted distance of path $P=v_{0},v_{1},..,v_{l}$, $\omega(P)$ to be $\omega(P):=\sum_{i=0}^{l-1}N\hat{x}_{v_{i}}$. For a subgraph $H$ of $G^{F}$ define the weighted distance $d_{\{\omega,H\}}(u,v)$ from $u$ to $v$ the minimum weight of a $u$-$v$ path in $H$. Define $d_{\omega}:=d_{\{\omega,G\}}$. For $U,W\subset V(G)$, define $d_{\omega,H}(U,W):=\min_{u\in U,w\in W}\ d_{\omega}(u,w)$. Define $d_{\omega}(U,W)=d_{\omega,G}(U,W)$. Define the weighted distance of a closed walk $P^{\prime}=v_{0},v_{1},..,v_{l}v_{0}$, $\omega(P^{\prime})$ to be $\omega(P^{\prime}):=\sum_{i=0}^{l}N\hat{x}_{v_{i}}$. The results in this paper could also be shown by instead defining the weight of each vertex to be $\hat{x}_{v}$ and instead defining the layers (see later) to be the vertices at distance a multiple of $1/N$ from a given set of vertices. Since $\hat{x}$ is feasible the following result holds.
Proposition 5.
For any $W\subset V(H)$ they define “layers” $L_{i}=\{v\in H:\ \ d_{H}(W,v)=i\}$ the nodes at distance $i$ from $W$. They show that the cost of all layers $L_{0},L_{1},...$ is $\sum_{v\in V(G)}N\hat{x}_{v}$. This is very useful for us as we will use this to show that one layer in $L_{1},...L_{m}$ has cost at most $\frac{1}{m}\sum_{v\in V}N\hat{x}_{v}$. However, the bypassing operation and replacing a node with a chain operation of [7] do not preserve the genus of the graph. We instead define the notion of weighted distance $d_{\omega}$. Denote the $i$-th layer from $W$ as $L_{i}:=\{v\in V:i\geq d_{\omega}(W,v)>i-\omega(v)\}$ the set of nodes for which the distance from $W$ to $v$ is at most $i$, but for which the distance plus the weight of $v$ is more than $i$. One can see that $v$ lies in $N\hat{x}_{v}$ different $L_{i}$, which is analogous to how $H$ defined in [7] contains $N\hat{x}_{v}$ copies of $v$ each lying in different layers as well. In particular, a node of weight 0 does not lie in any $L_{i}$, which is analogous to how a vertex of weight 0 is bypassed in [7].
Consider the embedding of $C^{1}$ on our surface. Given a subgraph $W$ of $G$, denote by $g(W)$ the subset of our surface occupied by a vertex or edge of $W$. We want to define a “small” neighbourhood around $g(C^{1})$, not containing any vertices outside $C^{1}$, which we divide up into “left” of $g(C^{1})$ and “right” of $g(C^{1})$, which we do using the following propositions. These are slightly informal statements of the exact propositions we require, the precise statements appear in section 5.
Proposition 6.
Proposition 7.
We defer the proofs of 6 and 7 for now. We apply 6 to $g(C^{1})$. Let $L,R$ be as in 6 so that each $g(e)$ for $e\in E(G)\backslash E(C^{1})$ is disjoint from at least one of $L,R$ and for each $e\in E(G\backslash C^{1})$, $g(e)$ is disjoint from both $L,R$.
Figure 1: $L$ and $R$ from 6 in yellow and red respectively curve $C^{\prime}$ depicted in black. The curve $h$ leaving $C^{\prime}$ from the left and entering from the right is depicted in dark green. The closed curve formed by $h$ and the subcurve of $C^{\prime}$ between $h(0)$ and $h(1)$ depicted in light green forms a non-facial closed curve.
Let $u^{\prime}_{i,1},u^{\prime}_{i,2},..,u^{\prime}_{i,l_{i}}$ be the out neighbours of $v_{i}$ such that the edges $u^{\prime}_{i,t^{\prime}}v_{i}$ reach $v_{i}$ from the left, that is, the arc obtained from reversing the arc $v_{i}u^{\prime}_{i,t^{\prime}}$ of our graph reaches $v_{i}$ from the left. Let $w^{\prime}_{i,1},w^{\prime}_{i,2},..,w^{\prime}_{i,z_{i}}$ be the in neighbours of $v_{i}$ such that the edges $w^{\prime}_{i,t^{\prime}}v_{i}$ reach $v_{i}$ from the right. Subdivide each edge $v_{i}u^{\prime}_{i,t}$ into a path $v_{i}u_{i,t}u^{\prime}_{i,t}$ and each edge $w^{\prime}_{j,t^{\prime}}v_{j}$ into a path $w^{\prime}_{j,t^{\prime}}w_{j,t^{\prime}}v_{j}$ and give the new vertices $w_{j,t^{\prime}},u_{i,t}$ infinite cost. There is a natural embedding of our new graph on our surface by placing each $u_{i,t}$ where the midpoint of the curve $g(v_{i}u^{\prime}_{i,t})$ was embedded and likewise for $w_{j,t^{\prime}}$. By abuse of notation, we continue to call our graph $G$ and define $\hat{x}_{u_{i,t}}=\hat{x}_{w_{j,t^{\prime}}}=0$ for all $u_{i,t}$, $w_{j,t^{\prime}}$. Denote $U:=\cup_{i=1}^{l^{\prime}}\{u_{i,1},u_{i,2},..,u_{i,l_{i}}\}$ and $W:=\cup_{i=1}^{l^{\prime}}\{w_{i,1},w_{i,2},..,w_{i,z_{i}}\}$. For $X\subset[l^{\prime}]$, denote $U_{X}:=\cup_{i\in X}\{u_{i,1},u_{i,2},..,u_{i,l_{i}}\}$, $V_{X}=\{v_{i}:\ \ i\in X\}$ and $W_{X}:=\cup_{i\in X}\{w_{i,1},w_{i,2},..,w_{i,z_{i}}\}$.
Claim 1.
If $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}})\leq\frac{1}{12}N$, then the LP gap of the natural LP ($\mbox{P}_{\text{DFVS}}$) for $G^{F}$ has integrality gap $O(g)$.
Proof.
Figure 2: Nodes of $U_{\tau_{+}}$ and $V_{\tau_{-}}$ shown in blue.
If $j\notin\tau_{-}$, then $w_{j,t^{\prime}}$ is not reachable from $U$ in $G^{F}\backslash S$. If $i\notin\tau_{+}$, then $W$ is not reachable from $u_{i,t}$ in $G^{F}\backslash S$. Thus, if either $j\notin\tau_{-}$ or $i\notin\tau_{+}$ then there is no path from $u_{i,t}$ to $w_{j,t^{\prime}}$ in $G^{F}\backslash S$. Thus, $j\in\tau_{-}$ and $i\in\tau_{+}$. As $K^{\prime}$ is strongly connected, this implies that $G^{F}\backslash S$ contains a path from $V_{\tau_{-}}$ to $V_{\tau_{+}}$, which is not possible.
Now suppose that $d_{\omega}(V_{\tau_{-}},V_{\tau_{+}})\leq\frac{1}{12}N$. Let $i\in\tau_{-}$ and $j\in\tau_{+}$ be such that $d_{\omega}(v_{i},v_{j})\leq\frac{1}{12}N$. Let $P_{1},P_{2},P_{3}$ be $u_{a,t}$-$v_{i}$, $u_{j,t^{\prime}}$-$v_{b}$ and $v_{i}$-$v_{j}$ paths of weight at most $\frac{1}{12}N$, with the second last vertices of $P_{1},P_{2}$ being in $W$, for some $a,b$. Such paths exist as $i\in\tau_{-}$ and $j\in\tau_{+}$. If $a=i$, then $P_{1}v_{i}u_{a,t}$ is a cycle for which $\sum_{v\in P_{1}v_{i}u_{a,t}}\hat{x}_{v}<1$ which is a contradiction. So $a\neq i$, likewise $b\neq j$.
We claim $\omega(C^{1}_{(v_{a},v_{i})}),\omega(C^{1}_{(v_{j},v_{b})})\leq\frac{1}{12}N$. Suppose for a contradiction that $\omega(C^{1}_{(v_{a},v_{i})})>\frac{1}{12}N$. Since $\omega(C^{1})=N$, this implies that $\omega(C^{1}_{(v_{i+1},v_{a-1})})<N-\frac{1}{12}N$. Then the cycle $P_{1}C^{1}_{(v_{i},v_{a})}v_{a}u_{a,t}$ satisfies $\sum_{v\in V(P_{1}C^{1}_{(v_{i},v_{a})}v_{a}u_{a,t})}\hat{x}_{v}<1-\frac{1}{12}+\frac{1}{12}=1$ which is a contradiction. Likewise, $\omega(v_{j},v_{b})\leq\frac{1}{12}N$.
Now suppose that $i=j$. Suppose for a contradiction that $C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{j},v_{b})}\neq\{v_{i}\}$. Let $v\in(C^{1}_{(v_{a},v_{i})}\cap C^{1}_{(v_{j},v_{b})})\backslash v_{i}$. Let $Q_{1}$ be a path from $v$ to $v_{i}$ in $C^{1}_{(v_{a},v_{i})}$ and $Q_{2}$ a path from $v_{i}$ to $v$ in $C^{1}_{(v_{i},v_{b})}$. Then $Q_{1}Q_{2}$ is a closed walk of weighted distance at most $\frac{1}{6}N$, which is a contradiction.
Claim 2.
Proof.
Refer to captionRefer to caption
Claim 3.
Proof.
Recall that any dicycle of $G^{F}\backslash R^{\prime}$ is contained in either $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or $G^{F}\backslash(P_{2}\cup C^{1}_{(v_{j+1},v_{b})})$. Since $v_{j}$ is not contained in any dicycle of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ it follows that any dicycle of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ is either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$. By 7, $g(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ and $g(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ are nonfacial.
Definition 3.
[1, 14] Given a simple closed curve $f$ on a surface without boundary $Q$, not dividing the surface into 2 regions, we say $Q^{\prime}$ is obtained by doing surgery along $f$ if $Q^{\prime}$ is obtained as follows. “Thicken” $f$ to obtain a cylinder and remove this cylinder from $Q$, call this resulting surface $Q^{\prime\prime}$. The boundary of $Q^{\prime\prime}$ consists of 2 circles we “glue” two cones $N_{1},N_{2}$ along these circles and call this final surface $Q^{\prime}$.